OpenStudy (anonymous):

exact value of sin 210° using the unit circle

OpenStudy (anonymous):

 so: sin(120) = sin(30) = -1/2 --------------------------------------------------------------------- OR 180-the angle° = the reference angle. so: 180 - 210 = -30 so: sin(-30) = -sin(30) = -(1/2) = -1/2

OpenStudy (anonymous):

Use the following guide 1. If the given angle is between 0 and 90 the the given angle is equal to the reference angle 2. If the given angle is between 90 and 180 the reference angle is equal to 180 - given angle 3. If the given angle is between 180 and 270 the reference angle is equal to the given angle - 180 4. If the given angle is between 270 and 360 the reference angle is equal to 360 - given angle Because the given angle is 210 the reference angle in this case is 210 - 180 = 30 The sign in this case is negative because is in the third quadrant Sin 210 = - sin (30) = -.5